OpenStudy (anonymous):

how can i find the ordered pairs of x^3 - y^3 = 721?

OpenStudy (dan815):

what are ordered pairs

OpenStudy (dan815):

u wanan know the values of x and y that work for this?

OpenStudy (anonymous):

yes, I want to know the work as to how i can get to them

OpenStudy (dan815):

i see

OpenStudy (dan815):

i dunno lol weird question, thers a bunch of ways for finding some ordered pairs

OpenStudy (dan815):

cant u just solve the equation for where y = 1 , y = 2 , y=3

OpenStudy (anonymous):

they have to be integers

OpenStudy (dan815):

well that will give u values of x too

OpenStudy (anonymous):

There are only 4 (x,y) pairs

OpenStudy (dan815):

i dunno agents giving u an answer tho

OpenStudy (agent0smith):

Hmm, you could start by using a difference of two cubes\[\Large x^3 - y^3 = (x-y)(x^2+xy+y^2) \] Since you want integers, you could find the factors of 721, since \[\Large (x-y)(x^2+xy+y^2) = 721\] means some integer times some integer equals 721. Factors of 721: 1, 7, 103, and 721 (I just googled them)

OpenStudy (agent0smith):

So we have \[\Large 1 \times 721 = 721\]\[\Large 7 \times 103 = 721\] \[\Large (x-y)(x^2+xy+y^2) = 721\] Let's first do 1*721... this gives \[\Large x- y = 1\]and \[\Large x^2+xy+y^2 = 721\] Or the other way around... \[\Large x- y = 721\]and \[\Large x^2+xy+y^2 = 1\] You can solve these since they're simultaneuous equations... oh and they also could both be negative... (-1)(-721)= 721...this gives you a couple more equations to solve (note they may not all have valid solutions).

OpenStudy (agent0smith):

Use substitution to solve those equations, and these ones. Now do it again for 7*103=721... giving \[\Large x- y = 7\]and \[\Large x^2+xy+y^2 = 103\] or \[\Large x- y = 103\]and \[\Large x^2+xy+y^2 = 7\] And again, (-7)(-103) = 721 so you may have to repeat them with negatives to find all solutions.

OpenStudy (dan815):

oh i see u want only integer solutions

OpenStudy (anonymous):

Thank you!

OpenStudy (agent0smith):

You're welcome. It'll be kinda tedious solving all those equations (there's really 8 total, if you check the negatives too, which you might need to). But it should give you your integer solutions.

OpenStudy (anonymous):

ok, thanks

OpenStudy (agent0smith):

Oh and I'm sure some of those might give non-integer solutions, which you can discard, or they might give you extraneous solutions.